OpenStudy (anonymous):

1.Evaluate n-11/13 for n=2/3 2.Solve y+22.461= -17.03 3.n+15.6/4=3.5

OpenStudy (anonymous):

2/3 = 26/39 11/13 = 33/39

OpenStudy (anonymous):

so it'd be 26/39 - 33/39 = -7/39

OpenStudy (anonymous):

y+22.461=-17.03 subtract 22.461 from both sides

OpenStudy (anonymous):

y= -39.491

OpenStudy (anonymous):

n+15.6/4=3.5 multiple both sides by 4/15.6

OpenStudy (anonymous):

n=.897 rounded